# How Propense Are Large Language Models at Producing Code Smells? A Benchmarking Study

### v1

@@ I Introduction
Large Language Models (LLMs) have been increasingly adopted to automate multiple tasks in software engineering (SE), including code completion [1, 2, 3], code summarization [4], program repair [5], clone detection [6], and test case generation [7]. The applications of LLMs in SE have evolved from classification tasks (e.g., defect detection, requirements triage, vulnerability severity) to generative tasks, such as sequence synthesis, where a prompt/context is given to instruct the LLMs to generate text (e.g., code summarization) and/or code (e.g., code generation) [8]. Ensuring the quality of the predictions made by those models is essential, regardless of the type of task. Generative tasks, in particular, require rigorous evaluation due to the complexity of structured information within the source code. Among the evaluation criteria, we have focused specifically on code smells.
@@
Code smells are defined as symptoms of poor design and implementation choices in SE. Previous studies have shown the negative impact of code smells on the comprehensibility and maintainability of code [9, 10]. Research has increasingly focused on studying the use of LLMs for detecting and refactoring code smells [11, 12, 13, 14, 9, 15, 16, 17, 18], and these models have demonstrated some success when their output is validated against the ground-truth. However, traditional metrics such as BLEU [19], CodeBLEU [20], ROUGE [21], and METEOR [22] fall short at assessing the LLMs’ tendency for introducing code smells while generating code. This limitation is particularly relevant for developers using commercial LLMs (e.g., ChatGPT, Copilot, and Claude  [^1] https://chatgpt.com, https://copilot.microsoft.com, https://claude.ai), which might generate code with smells like invalid-names, too-many-arguments, or unnecessary-lambda. Although LLMs are becoming more integrated into SE, and code smell detection has been extensively studied, there is still little known as to how likely LLMs are at generating smells.
@@
Despite an apparent success of using LLMs for automating software tasks, developers still face two key challenges: 1) they need more information to assess which LLM is more reliable, and 2) they cannot evaluate model performance beyond accuracy, as traditional metrics often overestimate code quality. This leads to important questions: What is the propensity of an LLM to generate code smells? and Which types of code smells are most propense to be generated?
@@
To address these concerns, we propose a new benchmark, CodeSmellEval, designed to estimate the propensity of an LLM at generating code smells. CodeSmellEval draws inspiration from Syntax Decomposition [23, 24] and introduces a model-agnostic evaluation metric called the Propensity Smelly Score (PSC). PSC analyzes the logits from the final layer of an LLM, providing insights into the model’s propensity of generating code smells. Additionally, our benchmark includes a new dataset, CodeSmellData, comprising $142k$ unique method-level code smells of $13$ different types, mined from GitHub.
@@
To demonstrate the utility of our benchmark, we devised an exploratory case study using CodeSmellEval in two LLMs (i.e., CodeLlama and Mistral). We assessed the distribution of PSC values for each type of code smells across both models. Our analysis revealed that both models are propense to generate the same types of code smells, with a few exceptions. For example, consider-merging-isinstance (R1701), chained-comparison (R1716), and broad-exception-caught (W0718) were among the top, whereas disallowed-name (C0104), too-many-arguments (R0913), and non-ascii-name (C2401) were bellow the PSC propensity threshold. We hope that our findings will shed light on the propensity of current LLMs to introduce code smells, enabling a more systematic and rigorous evaluation of code quality beyond canonical accuracy.
@@
To summarize, our key contributions are as the following: 1) a new metric for evaluating the propensity of LLMs to produce code smells during code generation (PSC), 2) a new dataset CodeSmellData of Python methods with Pylint code smells, 3) an exploratory case study that evaluates the propensity of two current popular open-source LLMs in producing code smells, and 4) and a set of notebooks and source code packages with instructions to replicate the experiments and use our benchmark [25].
@@ II-A Propensity Smelly Score (PSC)
Meaningful Structures $\mathbb{M}$. Using an alignment function ($\delta$), tokens $w_{i}\in w$ are grouped into a meaningful structure $\mu\in\mathbb{M}$ (Eq. 1). Then, an aggregation function ($\theta$) (Eq. 2) computes a central tendency statistic (e.g., median, mean, mode) of their probabilities, resulting in an overall probability estimate for predicting $\mu$ (i.e., propensity score). Once $P(w_{i}|w_{<i})$ is computed, the probability of each token’s predicted value is treated independently to calculate the overall estimate of $\mu$. Eq. 2 outlines how to compute the PSC for a meaningful concept $\mu$. Indexes that points out to smelly code are defined as $0\leq i\leq k\leq j\leq|w|$.
@@
Code Smells. The definition of the set of meaningful structures $\mathbb{M}$ used in function $\delta$ (Eq. 1) depends on the problem context. For example, token probabilities can be aggregated using syntax-based decomposition, based on elements defined by the grammar of a programming language (e.g., identifiers, conditionals, statements). In this paper, we propose defining the set of concepts $\mathbb{M}$ using types of method-level code smells. Table I illustrates the full taxonomy of code smells in $\mathbb{M}$ included in CodeSmellData.
Global Estimates. Finally, we compute the average of Eq. 2 for a given code smell type ($\mu$) across all code snippets $s$ in the dataset (CodeSmellData) to estimate the overall PSC of code smells generated by an LLM. We define a propensity threshold $\lambda=0.5$ to determine whether the LLM is propense to generate the code smell, with $\textit{PSC}\geq\lambda=0.5$ indicating a higher propensity. We use this estimate to answer RQ1. The selected propensity threshold was selected considering the findings of Karpaty et al. [26]. In addition, Fig. 2 depicts an example of computing PSC for a Python snippet containing the code smells: invalid-name with a PSC of $0.6$ above the propensity threshold and comparison-of-constants with a PSC of $0.206$.
@@ II-B CodeSmellData
*For detailed descriptions of each code smell, refer to the Pylint documentation [27].
To mitigate the risk of data leakage during the evaluation of code smells in generation tasks, we created a new dataset of code smells: CodeSmellData using Galeras [28]. We mined popular open-source Python repositories from GitHub, extracting a total of $232,715$ methods. The selection criteria included repositories published between January 2022 and December 2024, with a minimum of $3,500$ stars, ensuring that only well-maintained and highly rated repositories were considered. We then applied Pylint [27] analysis to these methods, detecting $79,574$ with at least one code smell.
Dataset Curation. From this filtered set of methods, we identified $156,151$ instances of code smells across $30$ distinct types. Using random sampling, the two authors manually validated instances from each code smell type to confirm true positives (TP) and false positives (FP) with an $80\%$ confidence level and $15\%$ margin error. We found that five code smells—inconsistent-return-statements, too-many-branches, too-many-return-statements, too-many-statements, and unbalanced-tuple-unpacking—had near to zero precision. For unbalanced-tuple-unpacking, the number of FPs exceeded the number of TPs, while for the other cases, Pylint failed to identify the exact location on the other code smell types of the smell within the method due to parsing errors. Additionally, $12$ code smells were excluded because they had fewer than $100$ instances, making them not representative (we observed that all PSC distributions tend to resemble a Gaussian distribution with at least $100$ instances). Finally, we ensured that all code smell instances in CodeSmellData originated from unique Python methods, avoiding any repetition of source code among code smell types. The resulting dataset consists of $142,817$ code smell instances from $13$ confirmed types, categorized into Refactoring, Warnings, and Conventions, as shown in Table I.
@@ III-A Results & Discussion
We believe that both LLMs are propense to generate the detected code smells, as they are trained on publicly available code, which often contains quality issues [29]. This assumption is supported by our dataset creation process, since we extracted code smells from mined code snippets in public repositories. However, a separate study is necessary to confirm this hypothesis, as this paper primarily focuses on demonstrating the use of CodeSmellEval. An in-depth evaluation of the underlying reasons why code smells appear in generated code is beyond the scope of this paper.
@@ V Future Plans
Based on the results of our case study, we demonstrated the computation of PSC for $13$ method-level code smells. Our next step is to conduct a systematic empirical evaluation that includes additional types of smells and a broader range of LLMs. This effort will involve mining more instances of underrepresented code smells in CodeSmellData. Specifically, we will analyze sampling error when computing PSC for each code smell to mitigate the risk of sampling bias and ensure the representativeness of all smell types in CodeSmellData. Furthermore, since our analysis has so far focused solely on method-level granularity, we plan to expand CodeSmellData to include higher-level smells, such as god-class and feature-envy. In future versions of CodeSmellData, we will use other static analysis tools alongside Pylint to reduce dependency on a single tool for identifying code smell locations, which is crucial for implementing the alignment function (Eq. 1) in CodeSmellEval.
Our proposed benchmark identifies the types of code smells that LLMs are propense to generate. Future research should focus on explaining this phenomenon by uncovering the reasons behind LLMs’ propensity to introduce specific types of code smells. For instance, by identifying the most relevant features in the input that influence the prediction of specific types of code smells. This research direction is essential for enhancing the trustworthiness[23] of LLMs and also developing defense techniques to mitigate such behavior. We believe interpretability techniques such as LIME [43] or SHAP [44] will help us to reach this goal.
VI Acknowledgments
This research has been supported in part by the NSF CCF-2311469, CNS-2132281, CCF-2007246, and CCF-1955853. We also acknowledge support from Cisco Systems. Any opinions, findings, and conclusions expressed herein are the authors’ and do not necessarily reflect those of the sponsors.

### v2

@@ I Introduction
Large Language Models (LLMs) have been used to automate multiple software engineering (SE) tasks, including code completion [1, 2, 3], code summarization [4], program repair [5], clone detection [6], and assertion generation [7]. LLMs’ applications have expanded from classification tasks (e.g., defect detection, requirements triage) to generative tasks, i.e., sequence synthesis in which LLMs generate text or code from a given prompt [8]. Ergo, ensuring LLMs’ prediction quality is crucial, particularly for SE tasks that require rigorous evaluation due to the complexity of source code. In this work, we evaluate the propensity of LLMs to generate or predict code smells.
@@
Code smells are symptoms of poor design and implementation choices, negatively affecting code comprehensibility and maintainability [9, 10]. Recent works have explored LLMs to detect and refactor smells [11, 12, 13, 14, 9, 15, 16, 17, 18], showing promising results when validated against ground truth. However, traditional metrics such as BLEU [19], CodeBLEU [20], ROUGE [21], and METEOR [22] fail to capture LLMstendency to introduce code smells at inference time. This tendency information is especially relevant for developers using commercial LLMs (e.g., ChatGPT, Copilot, and Claude [^1] https://chatgpt.com, https://copilot.microsoft.com, https://claude.ai), which may generate smelly snippets: invalid-names, too-many-arguments, or unnecessary-lambda. Although LLMs are largely investigated in SE life-cycle [23], little is known about the likelihood LLMs generate smelly code.
@@
Despite the apparent success of using LLMs for automating software tasks, developers still face two key challenges: 1) they need more information to assess which LLM is more reliable, and 2) they cannot evaluate model performance beyond accuracy, as traditional metrics often overestimate code quality. This leads to important questions: What is the propensity of an LLM to generate code smells? and Which types of code smells are most propense to be generated?
@@
To address these concerns, we propose a new benchmark, CodeSmellEval, designed to estimate the propensity of an LLM at generating smells. CodeSmellEval draws inspiration from Syntax Decomposition [24, 25] and introduces a model-agnostic evaluation metric called the Propensity Smelly Score (PSC). PSC analyzes the logits from the final layer of an LLM, providing insights into the model’s propensity of generating code smells. Additionally, our benchmark includes a new dataset, CodeSmellData, comprising $142k$ unique method-level code smells of $13$ different types, mined from GitHub.
@@
To demonstrate the utility of our benchmark, we designed an exploratory case study using CodeSmellEval in two LLMs (i.e., CodeLlama and Mistral). We assessed the distribution of PSC values for each type of smell in both models. Our analysis revealed that both models are propense to generate the same types of smell, with a few exceptions. For example, consider-merging-isinstance (R1701), chained-comparison (R1716), and broad-exception-caught (W0718) were among the top, whereas disallowed-name (C0104), too-many-arguments (R0913), and non-ascii-name (C2401) were bellow the PSC propensity threshold. We hope that our findings will shed light on the propensity of current LLMs to introduce code smells, enabling a more systematic and rigorous evaluation of code quality beyond canonical accuracy.
@@
To summarize, our key contributions are as follows: 1) a new metric to evaluate the propensity of LLMs to produce smells during code generation (PSC), 2) a new dataset CodeSmellData of Python methods with Pylint smells, 3) an exploratory case study to highlight the propensity of two current popular open-source LLMs to produce smells, and 4) and notebooks and code packages with instructions to replicate the experiments and use our benchmark [26].
@@ II-A Propensity Smelly Score (PSC)
Meaningful Structures $\mathbb{M}$. Using an alignment function ($\delta$), tokens $w_{i}\in w$ are grouped into a meaningful structure $\mu\in\mathbb{M}$ (Eq. 1). Then, an aggregation function ($\theta$) (Eq. 2) computes a central tendency statistic (e.g., median, mean, mode) of their probabilities, resulting in an overall probability estimate for predicting $\mu$ (i.e., propensity score). Once $P(w_{i}|w_{<i})$ is estimated, the likelihood of each expected token is used to calculate the value of $\mu$. Each token position probability is treated as independent because NTPs (Next Token Predictions) are generated auto-regressively.Eq. 2 outlines how to compute the PSC for a meaningful concept $\mu$. Indexes pointing out smelly code are defined as $0\leq i\leq k\leq j\leq|w|$.
@@
Code Smells. The definition of the set of meaningful structures $\mathbb{M}$ used in function $\delta$ (Eq. 1) depends on the problem context. For example, token probabilities can be aggregated using syntax-based decomposition, based on elements defined by the grammar of a programming language (e.g., identifiers, conditionals, statements). In this paper, we propose defining the set of concepts $\mathbb{M}$ using types of smells at method-level. Table I illustrates the full taxonomy of code smells in $\mathbb{M}$ included in CodeSmellData.
Global Estimates. Finally, we compute the average of Eq. 2 for a given code smell type ($\mu$) across all code snippets $s$ in the dataset (CodeSmellData) to estimate the overall PSC of code smells generated by an LLM. We define a propensity threshold $\lambda=0.5$ to determine whether the LLM is propense to generate the code smell, with $\textit{PSC}\geq\lambda=0.5$ indicating a higher propensity. We use this estimate to answer RQ1. The propensity threshold, based on the work of Karpaty et al. [27], is set at $0.5$ to indicate that code smells are more likely to be produced than flipping a coin. In addition, Fig. 2 depicts an example of computing PSC for a Python snippet containing the code smells: invalid-name with a PSC of $0.6$ above the propensity threshold and comparison-of-constants with a PSC of $0.206$.
@@ II-B CodeSmellData
*Refer to Pylint [28] for detailed smells’ description.
To mitigate the risk of data leakage during the smells evaluation, we created a new testbed: CodeSmellData using Galeras [29]. We mined popular open-source Python repositories from GitHub, extracting a total of $232,715$ methods. The selection criteria included repositories published between January 2022 and December 2024, with a minimum of $3,500$ stars, ensuring that only well-maintained and highly rated repositories were considered. We then applied Pylint [28] analysis to these methods, detecting $79,574$ with at least one code smell.
Dataset Curation. From this filtered set of methods, we identified $156,151$ instances of smells across $30$ distinct types. Using random sampling, two authors manually validated instances from each code smell type to confirm true positives (TP) and false positives (FP) with an $80\%$ confidence level and $15\%$ margin error. Discrepancies in the validation results were resolved through discussions among the authors. We found that five code smells—inconsistent-return-statements, too-many-branches, too-many-return-statements, too-many-statements, and unbalanced-tuple-unpacking—had near to zero precision. For unbalanced-tuple-unpacking, the number of FPs exceeded the number of TPs, while for the other cases, Pylint failed to identify the exact location on the other code smell types of the smell within the method due to parsing errors. Additionally, $12$ smells were excluded because they had fewer than $100$ instances, making them not representative (we observed that all PSC distributions tend to resemble a Gaussian distribution with at least $100$ instances). Finally, we ensured that all code smell instances in CodeSmellData originated from unique Python methods, avoiding any repetition of source code among code smell types. The resulting dataset consists of $142,817$ code smell instances from $13$ confirmed types, categorized into Refactoring, Warnings, and Conventions, as shown in Table I.
@@ III-A Results & Discussion
We believe that both LLMs are propense to generate detected code smells, as they are trained on publicly available code, which often contains quality issues [30]. Moreover, no evidence suggests that the training data of the selected LLMs has been curated to eliminate smells. This assumption is reinforced by our dataset creation process since we extracted smells from mined code snippets in public repositories. Nevertheless, a separate study is required to confirm this hypothesis as our goal is to demonstrate the utility of CodeSmellEval.
The results indicate that higher PSC values are correlated with a higher tendency to produce code smells, potentially reflecting models’ bias toward specific types of smells. However, since PSC is computed from conditioned generation (i.e., from code snippets containing a code smell), another study is needed to determine whether a correlation exists between PSC scores and the overall quality of the code generated by LLMs. While our observations highlight the need for larger-scale experiments and further validation, a detailed investigation into the underlying causes of code smells in generated code is beyond the scope of this paper.
@@ V Future Plans
Based on the results of our case study, we demonstrated the computation of PSC for $13$ method-level code smells. As a next step, we plan to conduct systematic experiments to evaluate the robustness of PSC by incorporating a wider variety of smells and a broader range of LLMs. This effort will involve mining more instances of underrepresented code smells in CodeSmellData. Specifically, we will analyze sampling error when computing PSC for each code smell to mitigate the risk of sampling bias and ensure the representativeness of all smell types in CodeSmellData. Furthermore, since our analysis has focused solely on method-level granularity, we plan to expand CodeSmellData to include higher-level smells, such as god-class and feature-envy. In future versions of CodeSmellData, we will use other static analysis tools alongside Pylint to reduce the dependency on a single tool for identifying smell locations, which is crucial to implement the alignment function (Eq. 1) in CodeSmellEval.
Our proposed benchmark identifies the types of code smells that LLMs are propense to generate. However, we acknowledge the need for empirical validation to demonstrate that PSC scores align with real-world developer expectations and impact. Future research should address how PSC supports practitioners in interpreting a model’s behavior and how it can be effectively used in practice. For instance, PSC could provide actionable insights by highlighting the types of smells a model is likely to introduce, enabling developers to assess model outputs more critically and prioritize mitigation strategies. Additionally, future work should uncover the reasons behind LLMs’ propensity to introduce specific types of smells, potentially by identifying the most relevant input features that influence the generation of these smells. This research direction is crucial for improving the trustworthiness [24] of LLMs and developing defense techniques to mitigate such behaviors. We believe interpretability techniques such as LIME [44] or SHAP [45] will be instrumental in achieving these goals.